# Length-Controlled AlpacaEval: A Simple Way to Debias Automatic Evaluators

### v1

@@ Abstract
LLM-based auto-annotators have become a key component of the LLM development process due to their cost-effectiveness and scalability compared to human-based evaluation. However, these auto-annotators can introduce complex biases that are hard to remove. Even simple, known confounders such as preference for longer outputs remains in existing automated evaluation metrics. We propose a simple regression analysis approach for controlling biases in auto-evaluations. As a real case study, we focus on reducing the length bias of AlpacaEval, a fast and affordable benchmark for chat LLMs that uses LLMs to estimate response quality. Despite being highly correlated with human preferences, AlpacaEval is known to favor models that generate longer outputs. We introduce a length-controlled AlpacaEval that aims to answer the counterfactual question: "What would the preference be if the model’s and baseline’s output had the same length?" To achieve this, we first fit a generalized linear model to predict the biased output of interest (auto-annotator preferences) based on the mediators we want to control for (length difference) and other relevant features. We then obtain length-controlled preferences by predicting preferences while conditioning the GLM with a zero difference in lengths. Length-controlling not only improves the robustness of the metric to manipulations in model verbosity, we also find that it increases the Spearman correlation with LMSYS’ Chatbot Arena from 0.94 to 0.98. We release the code and resulting leaderboard.
@@ 1 Introduction
Figure 1: Length-controlled AlpacaEval increases correlation with Chatbot Arena from 0.94 to 0.98. It is currently the benchmark with the highest correlation with Chatbot Arena.
Developing and improving NLP systems requires reliable, low-cost evaluations that can quantify progress. In closed-ended tasks, such as multiple-choice QA, such evaluations are straightforward to implement and trust (Novikova et al., 2017; Yeh et al., 2021). However, such evaluations cannot be applied to extremely open-ended settings such as instruction following for language models. Even neural reference-based evaluation metrics such as BERTscore (Zhang* et al., 2020) face challenges in those settings due to the difficulty of collecting a diverse set of references that can cover the space of valid outputs.
@@
Creating a way to debias automated evaluation metrics would be highly valuable – it would address the major drawback of LLM-based reference-free evaluations, and enable low-cost, accurate evaluations for developing NLP systems in many open-ended settings. Our work focuses on this challenge of taking an existing automated evaluation metric (e.g. AlpacaEval) and a suspected spurious correlate (e.g. length) and producing a debiased metric.
We propose a simple, interpretable debiasing strategy for automated evaluation metrics based on basic, regression-based adjustments for observational causal inference. We view spurious correlates – such as the length of the response – as undesirable mediators VanderWeele (2015) in a causal graph and use regression-based causal inference Hernán & Robins (2010) techniques to provide simple adjustments to automated evaluations that control for any suspected spurious correlation.
Applying this approach to the popular AlpacaEval benchmark, we show that controlling for length has significant positive effects on automated evaluation. We find that it is more correlated on average with LMSYS’ Chatbot Arena (Zheng et al., 2023) than both (length-uncontrolled) AlpacaEval and MT-bench, and that it is significantly more robust to gaming the evaluation by increasing the verbosity of the output.
@@
  • We propose a simple regression-based debiasing approach for automated evaluation that satisfies several desirable properties for an automatic evaluation metric.
  • We apply the approach to AlpacaEval, producing AlpacaEval-LC that is more robust to length-based spurious correlates.
@@ 2 Background and Problem Setting
Reference-free evaluation metrics have a long history, including classic methods (Louis & Nenkova, 2013) and more recent neural supervised learning methods (Kryscinski et al., 2020; Sinha et al., 2020; Goyal & Durrett, 2020). While this latter class of algorithms has become sufficiently accurate that they match inter-annotator agreement rates, other works have shown that such measurements are heavily confounded by spurious correlations such as perplexity and length (Durmus et al., 2022). Recently, there has been a push to leverage LLMs as a zero-shot, reference-free evaluation measure (Zheng et al., 2023; Dubois et al., 2023; Li et al., 2023; Lin et al., 2024). In the chatbot setting, two well-known such metrics are AlpacaEval and MT-bench, both of which query an LM (based upon GPT4) to attempt to assess the quality of a weaker LM’s output.
@@
AlpacaEval is an LLM-based automated evaluation metric – it operates on a fixed set of 805 instructions chosen to be representative of user interactions on the Alpaca web demo. For each instruction, both a baseline model $b$ (currently GPT-4 turbo) and the evaluated model $m$ produce responses. A GPT-4 turbo-based evaluator then compares the responses head-to-head and outputs the probability of preferring the evaluated model. A win rate is then computed as the expected probability that the auto-evaluator prefers the evaluated model’s output on the 805 instructions. This win rate serves as a performance measure of the evaluated LM chatbot.
Originally, AlpacaEval was designed as a development metric for the Alpaca chatbot (Taori et al., 2023) and AlpacaFarm simulator (Dubois et al., 2023). The metric was designed to control for certain biases, such as the order in which the model and baseline were presented, by randomizing their sequence. However, other factors like length and style effects were not controlled for, as the authors found that humans had similar biases on the analyzed data. Subsequent use of AlpacaEval as a leaderboard revealed that these uncontrolled biases could be significantly gamed by AI systems in ways that human biases couldn’t.
Important to our work is that AlpacaEval has several interpretable properties. As a win rate, its values are in $[0\%,100\%]$, it has a symmetry to baseline swaps, i.e, $\mathrm{AlpacaEval}(b,m)=100\%-\mathrm{AlpacaEval}(m,b)$, i.e., and comparing a baseline to itself is $\mathrm{AlpacaEval}(b,b)=50\%$ Any posthoc correction to AlpacaEval should maintain these properties, alongside the usual desiderata of being low-cost, accurate, and robust.
Chat Arena
The automated approach to pairwise evaluation in AlpacaEval can be viewed as a low-cost approximation to the Chatbot Arena Zheng et al. (2023), which aims to build real-world human evaluations through live interactions. The approach in the chatbot arena is that users are presented with a pair of anonymized language models, and they can send instructions to both language models simultaneously. The user receives responses from both LMs and rates the response that is higher quality. At the end, the head-to-head comparisons are converted to an Elo rating Elo & Sloan (1978) which serves as the model score. As a reminder, the difference of Elo ratings between two players can be converted to a win rate, and vis versa.
@@
In the remainder of this work, we will treat the Chatbot Arena as a silver standard that we wish to approximate. Although Chatbot Arena likely still contains biases (e.g. internet users may focus on surface features rather than “hard to measure” capabilities such as factuality), it represents the largest and most ecologically valid human evaluation process today.
Problem statement
We define the problem of pairwise evaluation of a language model in the following way. Given an instruction $x$ sampled from a distribution $p(x)$, a baseline model generates a response $z_{b}$ and the evaluated model generates a response $z_{m}$. A human annotator then produces a preference $y\in\{m,b\}$ indicating the model with the better response. An automated surrogate such as AlpacaEval is a (potentially randomized) predictor $f(z_{b},z_{m},x)$ that aims to approximate the corresponding human label $p(y=m|z_{b},z_{m},x)$. AlpacaEval’s current win rate is $\mathrm{winrate}(m,b)=100\cdot\mathbb{E}_{\scriptstyle x}\negthinspace\left[f(z_{b},z_{m},x)\right]$, i.e., the expected predicted preference of human annotators for the model response over the baseline’s response.
@@ 3 Length-Controlled AlpacaEval
A major challenge in building automated evaluations $f(z_{b},z_{m},x)$ is the spurious correlations problem. Specifically, consider a simple example in which there is a spurious correlate $c$ (e.g. length) such that heavily relying upon $c$ can be predictive of the human label $y$. The confounder $c$ is initially predictive of $y$ but becomes less predictive as model builders explicitly begin to optimize against the metric. Adopting this causal view, we ask our motivating question
What would the AlpacaEval metric be, if the outputs of all models had the same length as those of the baseline?
@@
Figure 2: Length-Controlled AlpacaEval predicts the direct effect of the model (green) on the auto-annotators’ preference (white) when controlling for undesirable mediators (red) and other useful features (gray).
@@
$\begin{split}q_{\theta,\phi,\psi}(y=m|z_{m},z_{b},x)&\vcentcolon=\\ \mathrm{logistic}&\bigg{(}\underbrace{\vphantom{\int}\theta_{m}-\theta_{b}}_{\text{\color[rgb]{0.13,0.55,0.13}Model}}+\underbrace{\vphantom{\int}\phi_{m,b}\cdot\mathrm{tanh}\left(\frac{\text{len}(z_{m})-\text{len}(z_{b})}{\text{std}(\text{len}(z_{m})-\text{len}(z_{b}))}\right)}_{\text{\color[rgb]{0.86,0.08,0.24}Length}}+\underbrace{\vphantom{\int}(\psi_{m}-\psi_{b})\gamma_{x}}_{\text{\color[rgb]{.5,.5,.5}Instruction}}\bigg{)}\end{split}$ (1)
The model and instruction terms are straightforward – they can be viewed as the log-linear contribution of the model ($m,b$) and each instruction’s difficulty ($\gamma$) on the baseline win rate. The length term is linear in a normalized length feature, where the normalizer standardizes the length to have unit variance and transforms this via a tanh, as differences in lengths should have strong diminishing returns on the log odds.
Importantly, this formula fulfills the identity property, i.e., $q(y=m\mid z_{b},z_{b},x)=0.5$ , and symmetry property, i.e, $q(y=m\mid z_{m},z_{b},x)=1.0-q(y=b\mid z_{b},z_{m},x)$ of the original win rate. Identity holds as the length term is zero due to having no difference in length, while the other two terms are zero as the coefficients are identical. For symmetry note that $\text{logistic}(x)=1-\text{logistic}(-x)$, and it is clear that swapping $m$ and $b$ flips the sign of the model and instruction terms. For the length term, the same is true as flipping $m$ and $b$ negates the length difference, and tanh is an odd function. More generally any additive term that is antisymmetric and centered around 0 would satisfy the desired properties.
Obtaining length corrected (LC) win-rate
@@
$\mathrm{winrate}^{LC}(m,b)=100\cdot\mathbb{E}_{\scriptstyle x}\negthinspace\left[\mathrm{logistic}\big{(}\theta_{m}-\theta_{b}+(\psi_{m}-\psi_{b})\gamma_{x}\big{)}\right].$ (2)
In other words, we simply remove the length term from the regression and compute the implied win rate.
@@
Training of the regression is simple and uses off-the-shelf libraries for fitting generalized linear models. Since our GLM uses a logit link function, we fit the model in Eq 1 using the cross-entropy loss $\mathcal{L}(\theta,\phi,\psi)=\mathbb{E}_{\scriptstyle p(y|z_{m},z_{b},x)p(z_{m},z_{b},x)}\negthinspace\left[q_{\theta,\phi,\psi}(y|z_{m},z_{b},x)\right]$.
In AlpacaEval’s leaderboard, we use a constant baseline $b$, so without loss of generality we can drop $\theta_{b},\psi_{b}$, which can be absorbed into the corresponding parameters for $m$. In total, for a leaderboard with $M$ models and $N$ instructions, our GLM contains $3M+N$ parameters to be estimated from $MN$ examples ($\theta_{m}$, $\phi_{m,b}$, $\psi_{m}$ for each model, $\gamma_{x}$ for each instruction). This will be overdetermined when $M$ and $N$ are both large, as in the case of AlpacaEval. However, to ensure our procedure is robust even for small $N$ and $M$, we use 5-fold cross-validation with $L_{2}$ regularization on weights to avoid potential overfitting.
@@
For the remaining regression coefficients, we simply fit $\theta$, $\phi$, and $\psi$ on the AlpacaEval predictions for each model separately, re-using the already estimated $\gamma_{x}$ as these do not depend on the model being evaluated. Fitting models separately is important as it implies that previously computed metrics won’t change when adding a new model to the leaderboard.
Finally, we have added an additional weak regularization on $\phi_{m,b}$ to prevent an adversary from performing attacks that intentionally truncate sequences that a model performs poorly on. In this case, the poor performance of the model would be perfectly correlated with the short length, and the model builder would be able to exploit the length corrections to boost the performance of the model. Adding a regularization term makes it so that any model performance issues would be explained by the model terms first, and then any residual effects would be captured by the length effects, as intended. The regularization is weak enough that we empirically found it to not affect non-adversarial models.

### v2

@@ Abstract
LLM-based auto-annotators have become a key component of the LLM development process due to their cost-effectiveness and scalability compared to human-based evaluation. However, these auto-annotators can introduce biases that are hard to remove. Even simple, known confounders such as preference for longer outputs remain in existing automated evaluation metrics. We propose a simple regression analysis approach for controlling biases in auto-evaluations. As a real case study, we focus on reducing the length bias of AlpacaEval, a fast and affordable benchmark for instruction-tuned LLMs that uses LLMs to estimate response quality. Despite being highly correlated with human preferences, AlpacaEval is known to favor models that generate longer outputs. We introduce a length-controlled AlpacaEval that aims to answer the counterfactual question: "What would the preference be if the model’s and baseline’s output had the same length?" To achieve this, we first fit a generalized linear model to predict the biased auto-annotator’s preferences based on the mediators we want to control for (length difference) and other relevant features. We then obtain length-controlled preferences by predicting preferences while conditioning the GLM with a zero difference in lengths. Length-controlling not only improves the robustness of the metric to manipulations in model verbosity, we also find that it increases the Spearman correlation with LMSYS Chatbot Arena from 0.94 to 0.98. We release the code and resulting leaderboard.
@@ 1 Introduction
Figure 1: AlpacaEval length-controlled increases correlation with the LMSYS Chatbot Arena from 0.94 to 0.98. It is currently the automatic benchmark with the highest correlation with Chatbot Arena.
Developing and improving NLP systems requires reliable, low-cost evaluations that can quantify progress. In closed-ended tasks, such as multiple-choice QA, such evaluations are straightforward to implement and trust (Novikova et al., 2017; Yeh et al., 2021). However, such evaluations cannot be applied to open-ended settings such as instruction following for language models. Even neural reference-based evaluation metrics such as BERTscore (Zhang* et al., 2020) face challenges in those settings due to the difficulty of collecting a diverse set of references that can cover the space of valid outputs.
@@
Creating a way to debias automated evaluation metrics would be highly valuable – it would address the major drawback of LLM-based reference-free evaluations and enable low-cost, accurate evaluations for developing NLP systems in open-ended settings. Our work focuses on this challenge of taking an existing automated evaluation metric (e.g. AlpacaEval) and a suspected spurious correlate (e.g. length) and producing a debiased metric.
We propose a simple, interpretable debiasing strategy for automated evaluation metrics based on basic, regression-based adjustments for observational causal inference. We view spurious correlates – such as the length of the response – as undesirable mediators VanderWeele (2015) in a causal graph and use regression-based causal inference Hernán & Robins (2010) techniques to provide simple adjustments to automated evaluations that control for any suspected spurious variables.
Applying this approach to the popular AlpacaEval benchmark, we show that controlling for length positively affects automated evaluation. We find that it is more correlated on average with the LMSYS Chatbot Arena (Zheng et al., 2023) than both (length-uncontrolled) AlpacaEval and MT-bench, and that it is significantly more robust to gaming the evaluation by increasing the verbosity of the output.
@@
  • We propose a simple regression-based approach to debias automated metrics, while maintaining desirable properties.
  • We apply the approach to AlpacaEval, producing AlpacaEval-LC, which is more robust to length-based spurious correlates.
@@ 2 Background and Problem Setting
Reference-free evaluation metrics, which aim to evaluate models on open-ended tasks without using any reference answer on that task, have a long history, including classic methods (Louis & Nenkova, 2013) and more recent neural supervised learning methods (Kryscinski et al., 2020; Sinha et al., 2020; Goyal & Durrett, 2020). While this latter class of algorithms has become sufficiently accurate that they match inter-annotator agreement rates, other works have shown that such measurements are heavily confounded by spurious correlations such as perplexity and length (Durmus et al., 2022). Recently, there has been a push to leverage LLMs as a zero-shot, reference-free evaluation measure (Zheng et al., 2023; Dubois et al., 2023; Li et al., 2023; Lin et al., 2024). In the chat setting, two well-known such benchmarks are AlpacaEval and MT-bench, both of which query an LM (based on GPT-4) to assess the quality of a weaker LM’s output.
@@
AlpacaEval is an LLM-based automated evaluation metric – it operates on a fixed set of 805 instructions chosen to be representative of user interactions on the Alpaca web demo (Taori et al., 2023). For each instruction, both a baseline model $b$ (currently GPT-4 turbo) and the evaluated model $m$ produce responses. A GPT-4 turbo-based evaluator then compares the responses head-to-head and outputs the probability of preferring the evaluated model. A win rate is then computed as the expected probability that the auto-evaluator prefers the evaluated model’s output on the 805 instructions. This win rate serves as a performance measure of the evaluated LM chatbot.
Originally, AlpacaEval was designed as a development metric for the Alpaca instruction-following model (Taori et al., 2023) and AlpacaFarm simulator (Dubois et al., 2023). The metric was designed to control for certain biases, such as the order in which the model and baseline were presented, by randomizing their sequence. However, other factors like length and style effects were not controlled for, as the authors found that humans had similar biases on the analyzed data. Subsequent use of AlpacaEval as a leaderboard revealed that these uncontrolled biases could be significantly gamed by AI systems in ways that human biases would likely not have.
Important to our work is that AlpacaEval has several interpretable properties. As a win rate, its values are in $[0\%,100\%]$, it has a symmetry to baseline swaps, i.e, $\mathrm{AlpacaEval}(b,m)=100\%-\mathrm{AlpacaEval}(m,b)$, i.e., and comparing a baseline to itself is $\mathrm{AlpacaEval}(b,b)=50\%$ Any post-hoc correction to AlpacaEval should maintain these properties, alongside the usual desiderata of being low-cost, accurate, and robustness.
Chatbot Arena
The automated approach to pairwise evaluation in AlpacaEval can be viewed as a low-cost approximation to the Chatbot Arena (Zheng et al., 2023), which aims to build real-world human evaluations through live interactions. The approach in Chatbot Arena is that users are presented with a pair of anonymized LMs, and they can send an instruction to both LMs simultaneously. The user receives responses from both LMs and rates the response that is of higher quality. In the end, the head-to-head comparisons are converted to Elo ratings (Elo, 1978) which serve as model scores. The difference in Elo ratings between two players can be converted to a win rate, and vice versa (Elo, 1978).
@@
In the remainder of this work, we will treat the Chatbot Arena as a silver standard, which we wish to approximate. Although Chatbot Arena likely still contains biases (e.g. internet users may focus on surface features rather than “hard to measure” capabilities such as factuality), it represents the largest publicly available human evaluation process today.
Setup
We define the problem of pairwise evaluation of a language model in the following way. Given an instruction $x$ sampled from a distribution $p(x)$, a baseline model generates a response $z_{b}$ and the evaluated model generates a response $z_{m}$. A human annotator then produces a preference $y\in\{0,1\}$ indicating whether the candidate response $z_{m}$ ($y=1$) is better than the baseline response $z_{b}$ ($y=0$). An automated surrogate such as AlpacaEval is a (potentially randomized) predictor $f(z_{m},z_{b},x)$ that aims to approximate the corresponding human label $p(y|z_{m},z_{b},x)$. AlpacaEval’s current metric is the win rate $\mathrm{winrate}(m,b)$ of the candidate model $m$ over the baseline $b$. I.e., the expected predicted preference of human annotators for the model response over the baseline’s response $\mathrm{winrate}(m,b)=100\cdot\mathbb{E}_{\scriptstyle x}\negthinspace\left[f(z_{m},z_{b},x)\right]$.
@@ 3 Length-Controlled AlpacaEval
A major challenge in estimating preferences $f(z_{m},z_{b},x)$ is the spurious correlations problem. Specifically, consider a simple example in which there is a spurious correlate $c$ (e.g. length) such that heavily relying upon $c$ can be predictive of the human preferences $y$ between the model’s and baseline’s output. The confounder $c$ is initially predictive of $y$ but could become less predictive as model builders explicitly begin to optimize against the metric. Adopting this causal view, we ask our motivating question:
What would the AlpacaEval win rate be, if the outputs of the evaluated model $\mathbf{m}$ had the same length as those of the baseline $\mathbf{b}$?
@@
Figure 2: Length-Controlled AlpacaEval predicts the direct effect of the model (blue) on the auto-annotators’ preference (white) when controlling for undesirable mediators (orange) and other useful features (gray).
@@
Specifically, for the first step, we will collect a list of ratings $\{x,z_{m},z_{b},m,b,y\}$ from the AlpacaEval leaderboard, and then fit a regression model $q_{\theta,\phi,\psi}$ to predict $y$ from $x,z_{m},z_{b},m,b$.
@@
$\begin{split}q_{\theta,\phi,\psi}(y=1|z_{m},z_{b},m,b,x)&\vcentcolon=\\ \mathrm{logistic}&\bigg{(}\underbrace{\vphantom{\int}\theta_{m}-\theta_{b}}_{\text{\color[rgb]{0.20703125,0.51953125,0.89453125}Model}}+\underbrace{\vphantom{\int}\phi_{m,b}\cdot\mathrm{tanh}\left(\frac{\text{len}(z_{m})-\text{len}(z_{b})}{\text{std}(\text{len}(z_{m})-\text{len}(z_{b}))}\right)}_{\text{\color[rgb]{1,0.62109375,0.42578125}Length}}+\underbrace{\vphantom{\int}(\psi_{m}-\psi_{b})\gamma_{x}}_{\text{\color[rgb]{.5,.5,.5}Instruction}}\bigg{)}.\end{split}$ (1)
The model and instruction terms are straightforward—they can be viewed as the log-linear contribution of the model ($m,b$) and each instruction’s difficulty ($\gamma$) on the baseline win rate. The length term is linear in a normalized length feature, where the normalizer standardizes the length to have unit variance and transforms this via a tanh, as differences in lengths should have strong diminishing returns on the log odds.
Importantly, this formula fulfills the identity property, i.e., $q(y=1\mid z_{b},z_{b},b,b,x)=0.5$ , and symmetry property, i.e, $q(y\mid z_{m},z_{b},b,m,x)=1.0-q(y\mid z_{b},z_{m},b,m,x)$ of the original win rate. Identity holds as the length term is zero due to having no difference in length, while the other two terms are zero as the coefficients are identical. For symmetry note that $\text{logistic}(x)=1-\text{logistic}(-x)$, and it is clear that swapping $m$ and $b$ flips the sign of the model and instruction terms. For the length term, the same is true as flipping $m$ and $b$ negates the length difference, and tanh is an odd function. More generally any additive term that is antisymmetric and centered around 0 would satisfy the desired properties.
Obtaining length corrected (LC) win rate
@@
$\mathrm{winrate}^{LC}(m,b)=100\cdot\mathbb{E}_{\scriptstyle x}\negthinspace\left[\mathrm{logistic}\big{(}\theta_{m}-\theta_{b}+(\psi_{m}-\psi_{b})\gamma_{x}\big{)}\right],$ (2)
i.e., we remove the length term from the regression and compute the implied win rate.
@@
Training of the regression is simple and uses off-the-shelf libraries for fitting generalized linear models. Since our GLM uses a logit link function, we fit the model in Eq 1 using the cross-entropy loss $\mathcal{L}(\theta,\phi,\psi)=\mathbb{E}_{\scriptstyle p(y|z_{m},z_{b},m,b,x)p(z_{m},z_{b},m,b,x)}\negthinspace\left[q_{\theta,\phi,\psi}(y|z_{m},z_{b},m,b,x)\right]$.
AlpacaEval’s leaderboard uses a constant baseline $b$, so without loss of generality we can drop $\theta_{b},\psi_{b}$, which can be absorbed into the corresponding parameters for $m$. In total, for a leaderboard with $M$ models and $N$ instructions, our GLM contains $3M+N$ parameters to be estimated from $MN$ examples ($\theta_{m}$, $\phi_{m,b}$, $\psi_{m}$ for each model, $\gamma_{x}$ for each instruction). This will be overdetermined when $M$ and $N$ are both large, as in the case of AlpacaEval ($M>128$, $N=805$). However, to ensure our procedure is robust even for small $N$ and $M$, we use 5-fold cross-validation with $L_{2}$ regularization on weights to avoid potential overfitting.
@@
For the remaining regression coefficients, we simply fit $\theta$, $\phi$, and $\psi$ on the AlpacaEval predictions for each model separately, re-using the already estimated $\gamma_{x}$ as these do not depend on the model being evaluated. Fitting models separately is important as it implies that previously computed metrics won’t change when adding a new model to the leaderboard. Of course, refitting $\gamma_{x}$ would give a better estimate of that parameter, but we don’t think it is worth constantly updating the leaderboard when a new model is added.
Finally, we added weak regularization on $\phi_{m,b}$ to prevent an adversary from performing attacks that intentionally truncate sequences that a model performs poorly on. In this case, the poor performance of the model would be perfectly correlated with the short length, and the model builder would be able to exploit the length corrections to boost the performance of the model. Adding a regularization term makes it so that any model performance issues would be explained by the model terms first, and then any residual effects would be captured by the length effects, as intended. The regularization is weak enough that we empirically found it to not affect non-adversarial models.